OpenStudy (tanya123):

1 + 1 +1 +1 + 1 1 + 1 + 1 + 1 +1 1 + 1 * 0 + 1 = ?

OpenStudy (inkyvoyd):

12

OpenStudy (hafeda):

u do the multiplication first, and after the addition

OpenStudy (tanya123):

@inkyvoyd that's wrong xD

OpenStudy (tanya123):

it's not 12

OpenStudy (sooobored):

its 12 if you follow order of operations PEMDAS or BODMAS -- doing the 0*1 first, then add the whole thing up if you punched it into a veryyyyyy simple calculator you'd get 1 since they work like 1+1 =ans ans+1= ans ans+1= ans ... ans*0= 0 0+1= 1 though if neither of those are right, then w/e

OpenStudy (tanya123):

Hm, no. That's not correct either :O

OpenStudy (hafeda):

LOL tanya can u solve it, because i really want to know it

OpenStudy (mrnood):

@tanya123 12 IS correct using the PEDMAS rules

OpenStudy (eliesaab):

There is no + sign at the end of the first two lines

OpenStudy (eliesaab):

If this is a trick question, you do not care about the first two lines, then 1+ 1*0 +1=2

OpenStudy (lolamona):

yea ama say its 2

imqwerty (imqwerty):

30

OpenStudy (mrnood):

ok - nice one :-) Teach us to read the question thoroughly!

OpenStudy (tanya123):

LOL